### v1

@@ 4.4.1 Model’s Behaviour on Clean Datasets
After addressing data quality issues and retraining the above models, we observed significant improvements in model behaviour and performance across all types of datasets: code-based, text-based, and metric-based. The impacts of data bugs were significantly diminished, and the models exhibited stable and normal behaviours during their training.
In code-based datasets, we found only 3.33% of models (1 out of 30) showing signs of inconsistent learning, compared to 83.37% previously. Model capacity improved substantially, with only 6.67% of models (2 out of 30) failing to capture complex patterns. Gradient stability also increased, with only 10.00% of models (3 out of 30) demonstrating instable gradients, down from 76.67% in the presence of data quality issues.
For text-based datasets, the chance of models having abnormal weight distribution decreased dramatically, with only 6.67% of models (2 out of 30) showing this behavior, compared to 93.33% before. Gradient skewness was largely prevented, with only 10.00% of models (3 out of 30) demonstrating skewed gradient distributions. Overfitting was significantly reduced, with only 6.67% of models (2 out of 30) showing this issue, down from 86.67% previously.
In metric-based datasets, sparse parameter updates were nearly prevented, with only 6.67% of models (2 out of 30) showing any signs of this issue. The vanishing gradient problem was significantly mitigated, with only 10.00% of models (3 out of 30) demonstrating this issue. Poor optimization was largely addressed, with only 13.33% of models (4 out of 30) showing this problem, compared to 73.33% previously.
The preprocessing issues, once addressed, also led to substantial improvements across all dataset types. In code-based datasets, only 6.67% of models (2 out of 30) showed slow convergence or incorrect learning patterns. For text-based data, only 10.00% of models (3 out of 30) demonstrated numerical instability or skewed bias distributions. In metric-based datasets, only 13.33% of models (4 out of 30) had exploding gradients or high variance in weight distribution.
4.5 Statistical Signifiance Tests
@@
We selected McNemar’s test pembury2020effective for the statistical tests due to our experimental design and paired nominal data, as it specifically evaluates changes in binary outcomes (presence/absence of symptoms) within the same models. While McNemar’s test establishes statistical significance to test our null hypothesis that data quality issues and model symptoms are independent, the odds ratio measures the strength of their dependence szumilas2010explaining . In our analysis, an odds ratio of 1 would support our null hypothesis, which indicates that there is no association between data quality issues and model symptoms. Conversely, odds ratios greater than 1 would support our alternative hypothesis, suggesting that the presence of data quality issues increases the likelihood of observed symptoms. Moreover, odds ratios below 1 would indicate that data quality issues decrease the likelihood of model symptoms.
McNemar’s test results consistently rejected the null hypothesis across all symptoms ($p<0.01$). The odds ratios, measuring the strength of association between data quality issues and symptoms, were substantially greater than 1 in all cases, ranging from $15.53$ for abnormal weight distribution to $806.10$ for gradient skewness, strongly supporting our alternative hypothesis. The most pronounced dependencies were observed in inconsistent learning, gradient instability, and overfitting ($p<0.0001$), indicating that these symptoms were exclusively present when data quality issues existed. Even the weakest association, found in the symptom ”high variance in weight distribution” ($p=0.0098$), demonstrated a statistically significant association between the symptom and data quality issues.
@@ 5 Implications for Stakeholders
Software Engineers: Software engineers can leverage our insights and implement continuous data monitoring systems that track properties such as weight distributions, gradient patterns, and learning stability across different layers. These systems can be used to automatically detect anomalies in training patterns and data distributions, enabling early identification of potential data quality issues before they are deployed in real-world systems. Our insights on the effects of concept drift can be used to establish processes for regular model retraining and deployment, ensuring their relevance over time. Moreover, the XAI techniques from our study can help software engineers debug and analyze model behaviour, identify potential biases, and address any issues before model deployment.
@@
Deep Learning Engineers: Deep learning engineers can leverage our empirical findings to establish automated data quality assurance, design validation mechanisms and integrate them into MLOps pipelines. The identified data quality issues and their manifestations can serve as the basis for designing effective validation rules, statistical tests, and domain-specific constraints. Furthermore, our study emphasizes the importance of robust preprocessing pipelines, which can guide deep learning engineers in building modular and extensible preprocessing components that adapt to different data types and evolving data characteristics. Our findings enable deep learning engineers to develop optimized training procedures with built-in validation mechanisms to ensure stable model convergence. Additionally, they can implement comprehensive monitoring and alerting systems that track critical indicators like weight distributions, gradient flows, and learning patterns, allowing rapid detection and response to potential data quality issues during model training.
@@ 8 Conclusion and Future Work
Deep learning systems suffer from data bugs that can significantly impact their reliability and trustworthiness. These bugs are particularly challenging to detect and resolve as they originate from various sources, including training data quality issues and preprocessing errors. In this study, we investigate the impact of data quality and preprocessing issues on the training behaviours of deep learning models targeting software engineering tasks. We target three prevalent data types: code-based, text-based, and metric-based data in the context of three tasks: vulnerability detection, duplicate bug report detection, and defect prediction. First, we replicated three state-of-the-art baselines, introducing data quality issues, omitting preprocessing operations and capturing their training-time behaviours using advanced tools (e.g., Weights & Biases). We then employed advanced explainable AI techniques to gain insights into the models’ learning processes and internal representations. Our key findings shed light on the impacts of data quality issues, such as label noise, concept drift, and class imbalance. These issues can lead to biased model behaviour, inconsistent learning across layers, reduced capacity to capture complex patterns, gradient instability during model training, overfitting to noisy patterns, vanishing gradients, poor optimization, and impaired generalization capabilities.
Additionally, the absence of necessary preprocessing operations can result in errors in parameter initialization, incorrect learning patterns, and biased feature extraction. This study emphasizes the critical importance of high-quality data and robust preprocessing pipelines for effective training of deep learning models in software engineering tasks. Our derived insights serve as a foundation for future work in developing data cleaning techniques, continuous data quality monitoring systems, and automated preprocessing pipelines tailored to different data types within the field of software engineering. Our findings can benefit debugging solutions by highlighting potential data-related issues that may impact model performance, developing more robust preprocessing techniques, and informing the design of automated quality assessment tools for machine learning pipelines in software engineering. These insights can also help practitioners identify and mitigate common data quality issues before negatively impacting model training and deployment.
Data Availability Statement (DAS)
@@
Conflict of Interest
The authors declare that they have no conflict of interest.

### v2

@@ 4.4.1 Model’s Behaviour on Clean Datasets
After addressing data quality and preprocessing issues and retraining the above models, we observed significant improvements in the behaviour of the deep learning models and their performance across all types of datasets: code-based, text-based, and metric-based. The impacts of data bugs were significantly diminished, and the models exhibited stable and normal behaviours during their training.
In code-based datasets, we found only $5.00\%$ of models showing signs of near-zero biases, compared to $81.67\%$. The quality of model weights improved substantially, with only $6.67\%$ showing the symptom of smaller weights. Gradient stability also increased, with only $8.33\%$ of models demonstrating unstable gradients, down from $68.33\%$ in the presence of data quality issues. For text-based datasets, the chance of models having abnormal weight distribution decreased dramatically, with only $6.67\%$ of models showing this behavior, compared to $85.00\%$ before. Gradient skewness was largely prevented, with only $8.33\%$ of models demonstrating skewed gradient distributions. Overfitting was significantly reduced, with only $5.00\%$ of models showing this issue, down from $85.00\%$ previously. In metric-based datasets, sparse parameter updates were nearly prevented, with only $6.67\%$ of models showing any signs of this issue. The vanishing gradient problem was significantly mitigated, with only $8.33\%$ of models demonstrating this issue. Poor optimization was largely addressed, with only $11.67\%$ of models showing this problem, compared to $70.00\%$ previously.
When models were trained with proper preprocessing techniques, we observed substantial improvements across all dataset types. In code-based datasets, we found only $3.33\%$ of models showing unstable convergence patterns, compared to $70.00\%$ previously. The quality of the bias distribution improved significantly, with only $5.00\%$ showing extreme bias values. For text-based datasets, the occurrence of extreme weights decreased dramatically, with only $5.00\%$ of models showing this behaviour, compared to $68.33\%$ before. Bias distribution issues were largely eliminated, with only $6.67\%$ of models demonstrating skewed distributions. In metric-based datasets, exploding gradients were nearly eliminated, with only $5.00\%$ of models showing signs of this issue. The high variance in weight distribution was significantly mitigated, with only $6.67\%$ of models demonstrating this issue.
4.4.2 Statistical Signifiance Tests
@@
We selected McNemar’s test pembury2020effective for the statistical tests due to our experimental design and paired nominal data, as it specifically evaluates changes in binary outcomes (presence/absence of symptoms) within the same models. While McNemar’s test establishes statistical significance to test our null hypothesis that data quality issues and model symptoms are independent, the odds ratio measures the strength of their dependence szumilas2010explaining . In our analysis, an odds ratio of 1 would support our null hypothesis, which indicates that there is no association between data quality issues and model symptoms. Conversely, odds ratios greater than one would support our alternative hypothesis, suggesting that the presence of data quality issues increases the likelihood of observed symptoms. Moreover, odds ratios below 1 would indicate that data quality issues decrease the likelihood of model symptoms.
McNemar’s test results consistently rejected the null hypothesis across all symptoms for data quality and data preprocessing issues ($p<0.01$). The odds ratios, measuring the strength of association between data quality issues and symptoms, were substantially greater than 1 in all cases, ranging from $15.53$ for abnormal weight distribution to $806.10$ for gradient skewness, strongly supporting our alternative hypothesis. The most pronounced dependencies were observed in near-zero biases, gradient instability, and overfitting ($p<0.0001$), indicating that these symptoms were exclusively present when data quality issues existed. Even the weakest association, found in the symptom ”high variance in weight distribution” ($p=0.0098$), demonstrated a statistically significant association between the symptom and data quality issues.
@@ 5.1 Implications for Stakeholders
Software Engineers: Software engineers can leverage our insights and implement continuous data monitoring systems that track properties such as weight distributions, gradient patterns, and learning stability across different layers of a model. These systems can be used to automatically detect anomalies in the training process and data distributions, enabling early identification of potential data quality issues before the models are deployed in real-world settings. Our insights on the effects of concept drift can be used to establish processes for regular model retraining and deployment, ensuring their relevance over time. Moreover, the XAI techniques from our study can help software engineers debug and analyze model behaviour, identify potential biases, and address any issues before their model deployment.
@@
Deep Learning Engineers: Deep learning engineers can leverage our empirical findings to establish automated data quality assurance and design validation mechanisms and integrate them into MLOps pipelines. The identified data quality issues and their manifestations can serve as the basis for designing effective validation rules, statistical tests, and domain-specific constraints. Furthermore, our study emphasizes the importance of robust preprocessing pipelines, which can guide deep learning engineers in building modular and extensible preprocessing components that adapt to different data types and evolving data characteristics. Our findings enable deep learning engineers to develop optimized training procedures with built-in validation mechanisms to ensure stable model convergence. Additionally, they can implement comprehensive monitoring and alerting systems that track critical indicators like weight distributions, gradient flows, and learning patterns, allowing rapid detection and response to potential data quality issues during model training.
5.2 Building Robust Data Pipelines for Deep Learning Model Training
Based on our findings, we recommend a systematic data pipeline for training models targeting software engineering tasks. It has four steps: data collection and versioning, quality validation, preprocessing, and monitoring throughout the training process.
First, we recommend establishing data collection and versioning protocols to promote reproducibility and traceability in the process. In this step, data should be collected from appropriate sources: code-based data from GitHub repositories and CI/CD logs, text-based data from bug trackers and commit messages, and metric-based data from static analysis tools. Organizations should implement provenance tracking through comprehensive metadata logging and data version control integration for systematic versioning. To maintain data integrity, raw data should be preserved in immutable formats such as Parquet files.
Second, we recommend implementing quality validation mechanisms for the captured data above. This phase can incorporate three key validation steps: label noise detection using confidence learning northcutt2021confident , concept drift monitoring through time-series analysis and drift detection algorithms, and class imbalance assessment using distribution analysis and adaptive sampling strategies.
Third, developing type-specific data preprocessing strategies for different data modalities is recommended. For text-based data, data should be extracted in chronological order from SE documentation, segmented into appropriate units, noisy content should be removed, and standard NLP preprocessing with tokenization should be applied. For code-based data, ASTs should be generated, code should be normalized and standardized, and token-based representations should be created. For metric-based data, duplicates should be eliminated, missing values should be handled, and feature normalization and scaling should be applied to ensure consistent ranges.
Finally, we recommend establishing Training Process Monitoring to implement comprehensive quality control during model training. The pipeline should continuously analyze parameter distributions through IQR and Z-score analyses for outlier detection or skewness and kurtosis measurements. Layer-wise distribution monitoring should track the behaviors of individual network components and compare them against the established thresholds. For weights and biases, both extremely small magnitudes (indicating potential vanishing gradients) and unusually large values (suggesting exploding gradients) should trigger monitoring alerts based on predefined thresholds. Based on observed anomalies, the model training can be interrupted and relevant details can be reported. The idea is to help subsequent analysis including root cause analysis, data repair, and re-training the model.
Based on our analysis across multiple datasets and model architectures, such a pipeline has the potential to help practitioners detect and address potential issues throughout the training process, ultimately enhancing model performance and robustness.
5.3 DataGuardian: Prototype for identifying data quality issues
To validate our findings and demonstrate their practical utility, we developed Data-Guardian replicationPackage , a monitoring system that can incorporate our parameter thresholds in a DL model for detecting data quality issues during model training. Our system continuously tracks parameter distributions, gradients, and layer-wise behaviours through a comprehensive set of statistical measures, including IQR, skewness, kurtosis, and distribution stability metrics.
DataGuardian offers flexibility in its monitoring approach through configurable thresholds. Users can either utilize our pre-defined thresholds, which were empirically determined through extensive experimentation, or define custom thresholds based on their specific requirements and domain knowledge. At the parameter level, it tracks weights and biases for symptoms such as abnormal weight distribution, overfitting, and skewed bias distributions. At the gradient level, Data-Guardian analyzes the flow of gradients through the network to detect problems such as gradient skewness, gradient instability, and vanishing or exploding gradients, which often indicate underlying data quality problems.
We evaluated DataGuardian’s effectiveness using multiple datasets across different domains: Juliet for code-based data, Mozilla for text-based data and JDT for metric-based data, using CodeBERT-based baseline models. They were not part of our threshold-related analysis. In experiments with buggy data, DataGuardian successfully identified training problems and their root causes using the our established thresholds in 12/15 runs for code-based data (Juliet), 11/15 for text-based data (Mozilla), and 12/15 for metric-based data (JDT). For scenarios with missing preprocessing steps, the system accurately detected training issues and identified the lack of preprocessing as the root cause in 10/15, 11/15, and 11/15 runs, respectively, for each dataset.
Our prototype also implements customizable interruption mechanism that can interrupt model training when certain pre-defined conditions are met. Users can also choose their thresholds concerning layer coverage of the encountered issues. This flexible approach helps distinguish between isolated anomalies (e.g., vanishing gradients) and systematic data quality problems while allowing users to balance between early detection and false positives. Upon detecting issues, Data-Guardian generates detailed diagnostic reports with specific metrics, potential root causes, and suggested mitigation strategies based on existing literature.
@@ 8 Conclusion and Future Work
Deep learning systems suffer from data bugs that can significantly impact their reliability and trustworthiness. These bugs are particularly challenging to detect and resolve as they originate from various sources, including training data quality issues and preprocessing errors. In this study, we investigate the impact of data quality and preprocessing issues on the training behaviours of deep learning models targeting software engineering tasks. We target three prevalent data types: code-based, text-based, and metric-based data in the context of three tasks: vulnerability detection, duplicate bug report detection, and defect prediction. First, we replicated three state-of-the-art baselines, introducing data quality issues, omitting preprocessing operations and capturing their training-time behaviours using advanced tools (e.g., Weights & Biases). We then employed advanced explainable AI techniques to gain insights into the models’ learning processes and internal representations. Our key findings shed light on the impacts of data quality issues, such as label noise, concept drift, and class imbalance. These issues can lead to near-zero biases, smaller weights, gradient instability during model training, slower convergence, overfitting to noisy patterns, vanishing gradients, higher loss, and impaired generalization capabilities. Future studies can reuse the same methodology and derive the impact of data quality and preprocessing issues on generation-related tasks such as code generation, code completion and code representation.
Additionally, the absence of necessary preprocessing operations can result in errors in extreme weights, skewed bias, and exploding gradients. This study emphasizes the critical importance of high-quality data and robust preprocessing pipelines for effective training of deep learning models in software engineering tasks. Our derived insights serve as a foundation for future work in developing data cleaning techniques, continuous data quality monitoring systems, and automated preprocessing pipelines tailored to different data types within the field of software engineering. Our findings can benefit debugging solutions by highlighting potential data-related issues that may impact model performance, developing more robust preprocessing techniques, and informing the design of automated quality assessment tools for machine learning pipelines in software engineering. These insights can also help practitioners identify and mitigate common data quality issues before negatively impacting model training and deployment.
Declarations
Funding and/or Conflicts of interests/Competing interests
This work was supported by the Natural Sciences and Engineering Research Council of Canada (Discovery Grant RGPIN‑03236). The authors also declare they have no conflict of interest.
Data Availability Statement
@@
Ethical Approval
Not applicable.
Informed Consent
Not applicable.
Author Contributions
  • Mehil B. Shah: Conceptualization, Methodology, Implementation, Investigation, Formal analysis, Visualization, Writing - original draft
  • Mohammad Masudur Rahman: Conceptualization, Supervision, Writing - review & editing, Validation, Conceptual guidance
  • Foutse Khomh: Supervision, Writing - review & editing, Validation, Conceptual guidance
Appendix A: Sample Output for Data Guardian
The Data Guardian system provides real-time monitoring of neural network training processes by detecting common training pathologies. As shown in Figure 6, the system identifies several critical issues that can significantly impact model performance:
Vanishing Gradients: This occurs when gradient values become extremely small during backpropagation, preventing effective weight updates in earlier network layers. The example shows this issue in two convolutional layers with gradient norms of 9.87e-08 and 3.45e-07, respectively. Without intervention, these layers would essentially stop learning.
Label Noise: Detected in the fc2.bias layer with a bias mean of 2.50e-03, this indicates potential inconsistencies in training labels. Label noise can lead to decreased generalization performance and longer convergence times as the model attempts to fit to contradictory patterns in the data.
Concept Drift: The embed_code.weight layer exhibits high weight variance (7.89e+00), suggesting that the statistical properties of the input data or target variables may be changing over time. This phenomenon often occurs in production systems where input distributions evolve beyond what was present in the original training data.
Exploding Gradients: At the opposite extreme of vanishing gradients, this issue manifests as excessively large gradient values—in this case, a gradient norm of 4.56e+02 in the fc1.weight layer. Exploding gradients can cause numerical instability, weight updates that overshoot optimal values, and in severe cases, NaN values that halt training entirely.
For each detected issue, Data Guardian automatically suggests appropriate remediation strategies, ranging from architectural adjustments to regularization techniques. These automated alerts allow machine learning engineers to intervene promptly during training, potentially saving significant computational resources and engineering time that would otherwise be spent debugging failed training runs.
Figure 6: Sample output from Data-Guardian